OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to x from the first equation is substituted into the second equation. x + y = 2 3x + 2y = 9

OpenStudy (anonymous):

3(2-y)+2y = 9

OpenStudy (lilg132):

x + y = 2 x = 2 - y 3(2-y) + 2y = 9